OpenStudy (anonymous):

. What is the solution of the system? Use substitution. 2x – y = –7 4x – y = –4 (1 point)(–1.5, 4) (1.5, 10) (4, –1.5) (–1.5, –2) What is the solution of the system? Use elimination. 5x + 7y = 32 8x + 6y = 46 (1 point)(8, 5) (1, 5) (7, 0) (5, 1) 16. Mike and Kim invest $12,000 in equipment to print yearbooks for schools. Each yearbook costs $5.00 to print and sells for $15.00. How many yearbooks must they sell before their business breaks even? (1 point)800 yearbooks 1,200 yearbooks 2,400 yearbooks 1,800 yearbooks 17. A baker has a bag of flour that is 11% wh

OpenStudy (cgreenwade2000):

DYODHDYODH

OpenStudy (texaschic101):

1. 2x - y = -7 --> -y = -7 - 2x --> y = 2x + 7 sub 2x + 7 in for y in the 2nd equation... 4x - y = -4 4x - (2x + 7) = -4 4x -2x - 7 = -4 2x = -4 + 7 2x = 3 x = 3/2 = 1.5 now sub 1.5 in for x in the 1st equation 2x - y = -7 2(1.5) - y = -7 3.0 - y = -7 -y = -7 - 3 -y = -10 y = 10 ANSWER : (1.5, 10) 2. 5x + 7y = 32 -->(8)5x + 7y = 32 8x + 6y = 46 -->(-5)8x + 6y = 46 ----------------- 40x + 56y = 256 (result of multiplying by 8) -40x - 30y = - 230 (result of multiplying by -5) ------------------add 26y = 26 y = 26/26 y = 1 now sub 1 in for y in either of the original equations... 5x + 7y = 32 5x + 7(1) = 32 5x + 7 = 32 5x = 32 - 7 5x = 25 x = 5 ANSWER : (5,1) 16. x stands for number of yearbooks sold 12,000 + 5(x) = 10(x) 12,000 = 10x - 5x 12,000 = 5x 12.000/5 = x 2400 ANSWER : 2400 yearbooks. I can't see the other question...